# Object-Centric Diffusion for Efficient Video Editing

### v1

@@ 4.2 Object-Centric 3D Token Merging
We further introduce an effective technique that promotes token merging in background areas while discouraging information loss in representations of foreground objects. More specifically, when applying ToMe, we associate each source token a binary value , obtained by aligning foreground masks to latent resolution, that specifies whether it belongs to the foreground region. Then, we simply account for in computing similarity between the source token and the destination token , as follows:
@@
where is a user-defined weighting factor such that corresponds to the original ToMe. By reducing the value of we reduce similarities of source tokens corresponding to edited objects, and therefore reducing their probability of being merged into destination tokens. The behavior of the weighted similarity metric can be appreciated in Fig. 4. As shown, the unmerged tokens (in blue) tend to stay more and more on the edited object while the merged tokens fall in background areas as decreases, although destination tokens are still sampled distributed across spatiotemporal volume.
Merging in 3D
@@
Diffusion-based video editing heavily relies on cross-frame attention operations to increase the temporal-consistency in generated frames [49]. Given its quadratic cost, this layer becomes the main computational bottleneck especially when increasing the number of frames it involves (Sec. 3). The standard ToMe only merges tokens in the spatial dimension. This is sub-optimal as it does not allow merging along the temporal dimension, which is the main source of redundancy for video. As a remedy, we apply ToMe within spatiotemporal volumes. Besides taking advantage of temporal redundancy, this strategy allows greater flexibility in choosing how to trade-off merging spatial vs. temporal information, by simply varying the size of the volumes.
@@ 5 Experiments
We apply our proposed strategy in the context of two families of video editing models. Specifically, we look into inversion-based video editing, where we rely on FateZero [33] as base model, and control-signal-based architectures, for which we optimize a ControlVideo [49] baseline. To provide fair comparisons in each setting, we rely on checkpoints and configurations provided by the corresponding authors. We evaluate each model using a benchmark of DAVIS [32] video sequences and edit prompts utilized in the original baseline methods. We refer the reader to supplementary materials for more-detailed setup.
As common in the literature and because we target zero-shot video editing, we optimize some of the hyperparameters (e.g. blending steps ratio , similarity re-weighting factor ) per-sequence for FateZero. For ControlVideo, we adopt the original hyperparameters reported in the paper.
Evaluation metrics:
@@ 5.1 Inversion-based Video Editing
Benchmark details
Following the benchmark in [33], we report a quantitative comparison based on 9 sequence-prompt pairs (7 from DAVIS and 2 in-the-wild). We include inversion-based video models such as Tune-A-Video + DDIM [46, 43] and TokenFlow + PnP [8, 44] in the evaluation. We also show results for a frame-based editing model, namely Framewise Null + p2p [31, 11]. We finally report results of the optimized baseline with fast sampling and ToMe (as discussed in Sec. 3.1 applied on top of FateZero), i.e., Optimized-FateZero.
@@
A qualitative comparison of state-of-the-art models is reported in Fig. 5, where we observe how our model enjoys the lowest latency, while ensuring a comparable editing quality. We report fidelity and latency measurements in Tab. 1, where our model achieves remarkable latency gains of 10 w.r.t. FateZero, and proves significantly faster than other state-of-the-art methods with comparable fidelity metrics. Although Optimized-FateZero is also a plausible model, we observe via visual assessment that its editing quality can be suboptimal, even though certain artifacts in generation might not be captured in CLIP fidelity metrics (for a clear example of this, please see the ablation in Fig. 7).
@@ 5.2 ControlNet-based Video Editing
Benchmark details
We follow the evaluation benchmark presented in [49] for evaluating ControlNet-based algorithms. The comparison is again based on sequences from DAVIS, and comprises 125 sequence-prompt pairs (detailed in the supplement), for which, per-frame depth maps are extracted using [36] and used as control signal. Besides our main baseline ControlVideo [49], we include Text2Video-Zero [17] in the comparisons. Once again, we will also report the performances of off-the-shelf optimizations as Optimized-ControlVideo.
@@
Fig. 6 illustrates some visual examples. As the figure shows, our proposal shows a significant speed-up over the baseline, obtaining a comparable generation quality. We also report a quantitative comparison of ControlNet-based methods in Tab. 2, observing similar outcomes. We notice that Text2Video-Zero is slightly faster than our method, due to its sparse instead of dense cross-frame attention. However, as we also observe, for the same reason it underperforms in terms of temporal-consistency among generated frames. We refer the reader to the supplementary material for more qualitative comparison.
Model | Tem-con | Cl-score | Latency (s) |
Text2Video-Zero [17] | 0.960 | 0.317 | 23.46 |
ControlVideo [49] | 0.972 | 0.318 | 152.64 |
Optimized-ControlVideo | 0.978 | 0.314 | 31.12 |
+ Object-Centric Diffusion | 0.977 | 0.313 | 25.21 |
@@ 5.3 Analysis
+ Object-Centric 3D ToMe | 0.967 | 9.05 | 10.54 |
@@
Ablation study
We ablate our main contributions within the FateZero base model. Specifically, we illustrate the impact of our proposals in Table 3, where we report Tem-con and latency as metrics, and in Fig. 7. Although off-the-shelf optimizations may grant reasonable speed-ups, we found that for extreme token reduction rates its editing results can suffer from generation artifacts. As the figure clearly shows, the addition of Object-Centric 3D ToMe can easily fix most defects on foreground regions by forcing ToMe to operate on other areas, and such benefit comes without major impact on latency. Finally, Object-Centric Sampling enables significant latency savings performing most diffusion iterations on foreground areas only. Somewhat surprisingly, we observe that Object-Centric Sampling helps the generation of background areas too: this is due to the fact that, as we run fewer denoising steps, its representations remain closer to the ones of the original video (resulting from inversion).
@@
The impact on latency of the proposed Object-Centric Sampling depends on the size of the foreground areas: in the presence of very small edits it can allow significant savings, whereas for prominent objects its editing cost would be only mildly-affected. We study this behavior in Tab. 4, where we divide 18 sequence-prompt pairs in three sets of 6 pairs each, depending on the number of tokens () within the object bounding box: Large with , Medium with and Small . For reference, the number of tokens of a whole frame is . As we show in Tab. 4, in the absence of Object-Centric Sampling, the latency of video editing does not depend on the size of the foreground objects. However, although we remark that our proposal reduces latency for all object sizes, we observe that gains are minor on large objects, as the foreground area covers most of the frame resolution and therefore Object-Centric Sampling is more akin to a regular diffusion process operating on the whole frame. Savings are more pronounced for medium and small sized objects, showing additional speed-ups of 1.7 and 2 respectively.
Object | #tokens | w/o Obj-Cen. Sampling | w/ Obj-Cen. Sampling | ||||
@@
size | () | Inv | Gen | Inv | Gen | ||
Large | 0.953 | 13.04 | 11.17 | 0.946 | 11.80 | 10.53 | |
Medium | 0.950 | 12.90 | 10.83 | 0.946 | 7.66 | 6.72 | |
Small | 0.979 | 12.94 | 10.88 | 0.975 | 6.30 | 5.85 |
Memory requirement for attention maps
As we noted that memory operations are a bottleneck for inversion-based architectures (see Sec. 3), we study the impact of our solutions on the memory required for attention operations. Specifically, we consider FateZero [33], whose attention maps sum up to 74.54 GB of floating point data (16bit). With simple off-the-shelf optimizations, this storage cost is reduced to 5.05 GB, and it can be further lowered to 4.22 GB by applying Object-Centric Sampling. This is a 17 reduction in memory footprint. This finding testifies a lower memory utilization corresponds to a reduced latency, and suggests that our solutions are effective in optimizing them.
@@ 6 Conclusion
In this paper, we introduced solutions for speeding up diffusion-based video editing. In this respect, we first presented an analysis of sources of latency in inversion-based models, and we identified and adopted some off-the-shelf techniques such as fast sampling and Token Merging that, when properly modified for the task, bring significant cost reduction without sacrificing generation quality. Furthermore, motivated by the fact that video editing typically requires modifications to specific objects, we introduce Object-Centric Diffusion, comprising techniques for i) isolating the merging of tokens in background regions only and ii) limiting most of the diffusion sampling steps on foreground areas. We validate our solutions on inversion-based and ControlNet-based models, where our strategies achieve 10 and 6 faster editing respectively.
Object-Centric Diffusion for Efficient Video Editing - Supplementary material
@@ A.1 Off-the-shelf optimizations of ToMe
Many inversion-based editing pipelines rely on sharing attention maps between inversion and generation stages (such as FateZero [33] and Plug-and-Play [44]). As such, when applying ToMe [4, 3], it is important that locations of destination (dst) and unmerged (unm) tokens are paired in the two stages, at the same attention layer and diffusion step. If that is not the case, tokens or attention maps coming from inversion are not compatible with the ones available at generation time. In practice, we compute which tokens to be merged during inversion, and use the same merges on generation attention maps when fusing. We found this strategy to be of primary importance, as testified by Fig. 3 (d-e) in the main paper.
@@
ToMe for stable diffusion [3] samples dst tokens randomly. When extending this to multiple frames, we initially sample the same random locations in each frame, finding this strategy to be sub-optimal. Instead, if we re-sample different random locations in each frame (or, in each temporal window in our 3D implementation), it allows us to preserve different information after merging. We found this to be beneficial, especially at higher merging rates (e.g. see Fig. 3 (e to f) in the main paper).
@@
In the original ToMe for stable diffusion [3], for each source (src) token, we search its corresponding match within a pool of dst tokens coming from the full spatial extent of the image (). The naive transposition of this strategy to our video use-case allows, for any src token, its candidate match to be searched within a pool of dst tokens coming from the full spatio-temporal extent of the video (). We find that this strategy for searching dst match, named hereby Temporally-Global Search, can lead to generation artifacts. Differently, we consider restricting the temporal extent of the dst pool to be the same temporal-window () as the src token, as in our Temporally-Windowed Search. As shown in Fig. A.8, the latter gives better reconstructions in general, whilst allowing more flexibility to control where merges are happening. This way, the user can also better trade-off the temporal-redundancy, smoothness and consistency by controlling the spatio-temporal window size.
@@
Sequence |
Source prompt |
Target prompts |
car-turn |
a silver jeep driving down a curvy road in the countryside. |
• a Porsche car driving down a curvy road in the countryside. • watercolor painting of a silver jeep driving down a curvy road in the countryside. |
train (in-the-wild)11footnotemark: 1 |
a train traveling down tracks next to a forest filled with trees and flowers and a man on the side of the track. |
• a train traveling down tracks next to a forest filled with trees and flowers and a man on the side of the track in Makoto Shinkai style. |
rabbit (in-the-wild)11footnotemark: 1 | a rabbit is eating a watermelon. | • pokemon cartoon of a rabbit is eating a watermelon. | |
Large object |
breakdance |
a man wearing brown tshirt and jeans doing a breakdance flare on gravel. |
• a woman with long-hair wearing green-sweater and jeans doing a breakdance flare on gravel. • a spiderman wearing red-blue spidersuit doing a breakdance flare on gravel. • a chimpanzee wearing a black jeans doing a breakdance flare on gravel. |
blackswan | a black swan with a red beak swimming in a river near a wall and bushes. | • a white duck with a yellow beak swimming in a river near a wall and bushes. • a pink flamingo with a red beak walking in a river near a wall and bushes. | |
car-turn |
a silver jeep driving down a curvy road in the countryside. |
• a Porsche car driving down a curvy road in the countryside. |
Medium object | lucia | a woman wearing a black dress with yellow handbag walking on a pavement. | • a woman wearing a white pant-suit with yellow handbag walking on a pavement. • a woman wearing a black dress and a hat with red handbag walking on a pavement. • a batman wearing a black bat-suit walking on a pavement. |
bear |
a brown bear walking on the rock against a wall. |
• a red tiger walking on the rock against a wall. • a yellow leopard walking on the rock against a wall. |
blackswan | a black swan with a red beak swimming in a river near a wall and bushes. | • a Swarovski crystal swan with a red beak swimming in a river near a wall and bushes. | |
Small object |
mallard |
a brown mallard running on grass land close to a lake. |
• a white duck running on grass land close to a lake. • a golden chicken running on grass land close to a lake. • a gray goose running on grass land close to a lake. |
soapbox |
two men driving a soapbox over a ramp. |
• two robots driving a mars-rover over a ramp. |
Sequence |
Source prompt |
Target prompts |
blackswan |
a black swan moving on the lake. |
• A black swan moving on the lake • A white swan moving on the lake. • A white swan moving on the lake, cartoon style. • A crochet black swan swims in a pond with rocks and vegetation. • A yellow duck moving on the river, anime style. |
boat |
a boat moves in the river. |
• A sleek boat glides effortlessly through the shimmering river, van gogh style. • A majestic boat sails gracefully down the winding river. • A colorful boat drifts leisurely along the peaceful river. • A speedy boat races furiously across the raging river. • A rustic boat bobs gently on the calm and tranquil river. |
breakdance-flare |
a man dances on the road. |
• A young man elegantly dances on the deserted road under the starry night sky. • The handsome man dances enthusiastically on the bumpy dirt road, kicking up dust as he moves. • A man gracefully dances on the winding road, surrounded by the picturesque mountain scenery. • The athletic man dances energetically on the long and straight road, his sweat glistening under the bright sun. • The talented man dances flawlessly on the busy city road, attracting a crowd of mesmerized onlookers. |
bus |
a bus moves on the street. |
• A big red bus swiftly maneuvers through the crowded city streets. • A sleek silver bus gracefully glides down the busy urban avenue. • A colorful double-decker bus boldly navigates through the bustling downtown district. • A vintage yellow bus leisurely rolls down the narrow suburban road. • A modern electric bus silently travels along the winding coastal highway. |
camel |
a camel walks on the desert. |
• A majestic camel gracefully strides across the scorching desert sands. • A lone camel strolls leisurely through the vast, arid expanse of the desert. • A humpbacked camel plods methodically across the barren and unforgiving desert terrain. • A magnificent camel marches stoically through the seemingly endless desert wilderness. • A weathered camel saunters across the sun-scorched dunes of the desert, its gaze fixed on the horizon. |
car-roundabout |
a jeep turns on a road. |
• A shiny red jeep smoothly turns on a narrow, winding road in the mountains. • A rusty old jeep suddenly turns on a bumpy, unpaved road in the countryside. • A sleek black jeep swiftly turns on a deserted, dusty road in the desert. • A modified green jeep expertly turns on a steep, rocky road in the forest. • A gigantic yellow jeep slowly turns on a wide, smooth road in the city. |
car-shadow |
a car moves to a building. |
• A sleek black car swiftly glides towards a towering skyscraper. • A shiny silver vehicle gracefully maneuvers towards a modern glass building. • A vintage red car leisurely drives towards an abandoned brick edifice. • A luxurious white car elegantly approaches a stately colonial mansion. • A rusty blue car slowly crawls towards a dilapidated concrete structure. |
car-turn |
a jeep on a forest road. |
• A shiny silver jeep was maneuvering through the dense forest, kicking up dirt and leaves in its wake. • A dusty old jeep was making its way down the winding forest road, creaking and groaning with each bump and turn. • A sleek black jeep was speeding along the narrow forest road, dodging trees and rocks with ease. • A massive green jeep was lumbering down the rugged forest road, its powerful engine growling as it tackled the steep incline. • A rusty red jeep was bouncing along the bumpy forest road, its tires kicking up mud and gravel as it went. |
cows |
a cow walks on the grass. |
• A spotted cow leisurely grazes on the lush, emerald-green grass. • A contented cow ambles across the dewy, verdant pasture. • A brown cow serenely strolls through the sun-kissed, rolling hills. • A beautiful cow saunters through the vibrant, blooming meadow. • A gentle cow leisurely walks on the soft, velvety green grass, enjoying the warm sunshine. |
dog |
a dog walks on the ground. |
• A fluffy brown dog leisurely strolls on the grassy field. • A scruffy little dog energetically trots on the sandy beach. • A majestic black dog gracefully paces on the polished marble floor. • A playful spotted dog joyfully skips on the leaf-covered path. • A curious golden dog curiously wanders on the rocky mountain trail. |
elephant |
an elephant walks on the ground. |
• A massive elephant strides gracefully across the dusty savannah. • A majestic elephant strolls leisurely along the lush green fields. • A mighty elephant marches steadily through the rugged terrain. • A gentle elephant ambles peacefully through the tranquil forest. • A regal elephant parades elegantly down the bustling city street. |
Sequence |
Source prompt |
Target prompts |
flamingo |
a flamingo wanders in the water. |
• A graceful pink flamingo leisurely wanders in the cool and refreshing water, its slender legs elegantly stepping on the soft sand. • A vibrant flamingo casually wanders in the clear and sparkling water, its majestic wings spread wide in the sunshine. • A charming flamingo gracefully wanders in the calm and serene water, its delicate neck curving into an elegant shape. • A stunning flamingo leisurely wanders in the turquoise and tranquil water, its radiant pink feathers reflecting the shimmering light. • A magnificent flamingo elegantly wanders in the sparkling and crystal-clear water, its striking plumage shining brightly in the sun. |
gold-fish |
golden fishers swim in the water. |
• Majestic golden fishers glide gracefully in the crystal-clear waters. • Brilliant golden fishers swim serenely in the shimmering blue depths. • Glittering golden fishers dance playfully in the glistening aquamarine waves. • Gleaming golden fishers float leisurely in the peaceful turquoise pools. • Radiant golden fishers meander lazily in the tranquil emerald streams. |
hike |
a man hikes on a mountain. |
• A rugged man is trekking up a steep and rocky mountain trail. • A fit man is leisurely hiking through a lush and verdant forest. • A daring man is scaling a treacherous and jagged peak in the alpine wilderness. • A seasoned man is exploring a remote and rugged canyon deep in the desert. • A determined man is trudging up a snowy and icy mountain slope, braving the biting cold and fierce winds. |
hockey |
a player is playing hockey on the ground. |
• A skilled player is furiously playing ice hockey on the smooth, glistening rink. • A young, agile player is energetically playing field hockey on the lush, green grass. • An experienced player is gracefully playing roller hockey on the sleek, polished pavement. • A determined player is passionately playing street hockey on the gritty, urban asphalt. • A talented player is confidently playing air hockey on the fast-paced, neon-lit table. |
kite-surf |
a man is surfing on the sea. |
• A muscular man is expertly surfing the gigantic waves of the Pacific Ocean. • A handsome man is gracefully surfing on the crystal clear waters of the Caribbean Sea. • A daring man is fearlessly surfing through the dangerous, choppy waters of the Atlantic Ocean. • An athletic man is skillfully surfing on the wild and untamed waves of the Indian Ocean. • A young man is confidently surfing on the smooth, peaceful waters of a serene lake. |
lab-coat |
three women stands on the lawn. |
• Three stunning women are standing elegantly on the lush green lawn, chatting and laughing. • Three young and vibrant women are standing proudly on the well-manicured lawn, enjoying the sunshine. • Three fashionable women in colorful dresses are standing gracefully on the emerald green lawn, taking selfies. • Three confident women with radiant smiles are standing tall on the soft, green lawn, enjoying the fresh air. • Three beautiful women, each dressed in their own unique style, are standing on the lush and verdant lawn, admiring the scenery. |
longboard |
a man is playing skateboard on the alley. |
• A young man is skillfully skateboarding on the busy city street, weaving in and out of the crowds with ease. • An experienced skateboarder is fearlessly gliding down a steep, curvy road on his board, executing impressive tricks along the way. • A daring skater is performing gravity-defying flips and spins on his board, effortlessly navigating through a challenging skatepark course. • A talented skateboarder is carving up the smooth pavement of an empty parking lot, creating beautiful patterns with his board and body. • A passionate skater is practicing his moves on a quiet neighborhood street, with the sound of his board echoing through the peaceful surroundings. |
mallard-water |
a mallard swims on the water. |
• A vibrant mallard glides gracefully on the shimmering water. • A beautiful mallard paddles through the calm, blue water. • A majestic mallard swims elegantly on the tranquil lake. • A striking mallard floats effortlessly on the sparkling pond. • A colorful mallard glides smoothly on the rippling surface of the water. |
mbike-trick |
a man riding motorbike. |
• A young man riding a sleek, black motorbike through the winding mountain roads. • An experienced man effortlessly riding a powerful, red motorbike on the open highway. • A daring man performing gravity-defying stunts on a high-speed, blue motorbike in an empty parking lot. • A confident man cruising on a vintage, yellow motorbike along the picturesque coastal roads. • A rugged man maneuvering a heavy, dusty motorbike through the rugged terrain of a desert. |
Sequence |
Source prompt |
Target prompts |
rhino |
a rhino walks on the rocks. |
• A massive rhino strides confidently across the jagged rocks. • A majestic rhino gracefully navigates the rugged terrain of the rocky landscape. • A powerful rhino marches steadily over the rough and rocky ground. • A colossal rhino plods steadily through the craggy rocks, undeterred by the challenging terrain. • A sturdy rhino confidently traverses the treacherous rocks with ease. |
surf |
a sailing boat moves on the sea. |
• A graceful sailing boat glides smoothly over the tranquil sea. • A sleek sailing boat cuts through the shimmering sea with ease. • A majestic sailing boat cruises along the vast, azure sea. • A vintage sailing boat bobs gently on the calm, turquoise sea. • A speedy sailing boat races across the glistening, open sea. |
swing |
a girl is playing on the swings. |
• A young girl with pigtails is joyfully swinging on the colorful swings in the playground. • The little girl, giggling uncontrollably, is happily playing on the old-fashioned wooden swings. • A blonde girl with a big smile on her face is energetically playing on the swings in the park. • The girl, wearing a flowery dress, is gracefully swaying back and forth on the swings, enjoying the warm breeze. • A cute little girl, dressed in a red coat, is playfully swinging on the swings, her hair flying in the wind. |
tennis |
a man is playing tennis. |
• The skilled man is effortlessly playing tennis on the court. • A focused man is gracefully playing a game of tennis. • A fit and agile man is playing tennis with precision and finesse. • A competitive man is relentlessly playing tennis with his opponent. • The enthusiastic man is eagerly playing a game of tennis, sweat pouring down his face. |
walking |
a selfie of walking man. |
• A stylish young man takes a selfie while strutting confidently down the busy city street. • An energetic man captures a selfie mid-walk, showcasing his adventurous spirit. • A happy-go-lucky man snaps a selfie as he leisurely strolls through the park, enjoying the sunny day. • A determined man takes a selfie while briskly walking towards his destination, never breaking stride. • A carefree man captures a selfie while wandering aimlessly through the vibrant cityscape, taking in all the sights and sounds. |
@@
As observed in [3], the high resolution UNet [39] stages are the most expensive in terms of self-attention (or, cross-frame attention) modules, and the ones that can benefit the most by applying ToMe. Contrarily to the original formulation which does not optimize low-resolution layers, we do apply ToMe in all layers as we observe it has a meaningful impact on latency. We however cap the minimum #tokens in low-resolution layers in order to avoid degenerate bottlenecks (e.g. collapsing to a single representation). Specifically, we maintain at-least 4 and 16 tokens per-frame after merging at and resolutions, respectively.

### v2

@@ 4.2 Object-Centric Token Merging
We further introduce an effective technique that promotes token merging in background areas while discouraging information loss in representations of foreground objects. More specifically, whenever applying ToMe, we associate each source token a binary value , obtained by aligning foreground masks to latent resolution, that specifies whether it belongs to the foreground region. Then, we simply account for in computing similarity between the source token and the destination token , as follows:
@@
where is a user-defined weighting factor ( corresponds to the original ToMe). By reducing the value of , we deliberately weaken the similarities of source tokens corresponding to edited objects, therefore reducing their probability of being merged. The behavior of the weighting factor can be appreciated in Fig. 4: as decreases, the unmerged tokens (in blue) tend to locate more and more on the edited objects, avoiding information loss in their locations.
Merging spatio-temporal token volumes
@@
Diffusion-based video editing heavily relies on cross-frame attention to increase the temporal-consistency in generated frames [47, 21], incurring in severe computational bottlenecks. The standard ToMe merges tokens only in the spatial dimension, underexploiting the most prominent redundancies in videos, which occur along the time axis. As a remedy, we apply ToMe within spatiotemporal volumes. This strategy takes advantage of temporal redundancy and allows flexibility in choosing how to trade-off merging spatial vs. temporal information by simply varying the size of the volumes.
@@ 5 Experiments
We test OCD in the context of two families of video editing models.
Specifically, we look into inversion-based video editing, where we rely on FateZero [32] as base model, and control-signal-based architectures, for which we optimize a ControlVideo [47] baseline.
To provide fair comparisons in each setting, we rely on checkpoints and configurations provided by the corresponding authors.
We evaluate each model using a benchmark composed of DAVIS [31] video sequences and edit prompts utilized in the original baseline methods.
Implementation details
For both video editing models, we pad the mask to a rectangular shape and implement the gather operation in Algorithm 1 as a crop around the foreground object. Although potentially suboptimal for non-rectangular objects, this strategy proves effective in practice (Tab. 2) and enables the use of regular dense convolutions within UNet, which are typically highly optimized and even faster than sparse counterparts. Such an approximation would however not be necessary for other convolution-free editing pipelines, e.g. fully based on transformer architectures [30]. As common in the literature and because we target zero-shot video editing, we optimize some of the hyperparameters (e.g. blending steps ratio , similarity re-weighting factor ) per-sequence for FateZero. For ControlVideo, we adopt the original hyperparameters reported in the paper. is selected in based on heuristics: a rule-of-thumb, is to use a lower value (preserving more foreground), when the object has complex motion or texture (i.e., it is challenging to reconstruct). As for the saliency masks, we utilize the segmentation masks available within the DAVIS dataset. We refer the reader to supplementary materials for more-detailed setup.
Evaluation metrics
@@ 5.1 Inversion-based Video Editing
Following the benchmark in [32], we report a quantitative comparison based on 9 sequence-prompt pairs. We include inversion-based video models such as Tune-A-Video + DDIM [44, 42], TokenFlow + PnP [7, 43] and VidToMe [21] in the evaluation. We also show results for a frame-based editing model, namely Framewise Null + p2p [29, 9]. We finally report results of the Optimized-FateZero baseline, obtained by the off-the-shelf accelerations in Sec. 3.1.
@@
A qualitative comparison of OCD and state-of-the-art models is reported in Fig. 5, where our model enjoys the lowest latency, while ensuring a comparable editing quality. We report fidelity and latency measurements in Tab. 1, where OCD achieves remarkable latency gains of 10 w.r.t. FateZero, and proves significantly faster than other state-of-the-art methods with comparable fidelity metrics. Although Optimized-FateZero is also very fast, we observe via visual assessment that its editing quality can be suboptimal, even though highly localized geneation artifacts might be overlooked by fidelity metrics (for clear examples of this, we refer to the ablation in Fig. 7).
@@ 5.2 ControlNet-based Video Editing
We follow the evaluation benchmark in [47] for evaluating ControlNet-based algorithms. The comparison comprises 125 sequence-prompt pairs from DAVIS (detailed in the supplement), for which per-frame depth maps are extracted using [35] and used as control signal. Besides our main baseline ControlVideo [47], we include Text2Video-Zero [15] in the comparisons. Once again, we will also report the performances of off-the-shelf optimizations (Optimized-ControlVideo).
@@
Fig. 6 illustrates some visual examples. As the figure shows, our proposal shows a significant speed-up over the baseline, obtaining a comparable generation quality. We also report a quantitative comparison of ControlNet-based methods in Fig. 6, observing similar outcomes. We notice that Text2Video-Zero is slightly faster than our method, due to its sparse instead of dense cross-frame attention. However, for the same reason it underperforms in terms of temporal-consistency among generated frames. We refer the reader to the supplementary material for additional qualitative comparison.
[…]
[table]
@@ 5.3 Analysis
| + Object-Centric ToMe | 0.967 | 9.05 | 10.54 |
@@
[…]
[table]
We ablate our main contributions within the FateZero base model. Specifically, we illustrate the impact of our proposals in Table 7, where we report Tem-con and latency as metrics, and in Fig. 7. Although off-the-shelf optimizations may grant reasonable speed-ups, we found that for extreme token reduction rates their editing results can suffer from generation artifacts. As the figure shows, the addition of Object-Centric ToMe can easily fix most defects on foreground regions by forcing ToMe to operate on other areas, a benefit that comes without major impacts on latency. Finally, Object-Centric Sampling enables significant latency savings performing most diffusion iterations on foreground areas only. Somewhat surprisingly, we observe that Object-Centric Sampling helps the generation of background areas too: this is due to the fact that, as we run fewer denoising steps, its representations remain closer to the ones of the original video (resulting from inversion).
| Saliency mask | Tem-con | Cl-score |
| GT mask | 0.967 | 0.331 |
| Grounded SAM [36] | 0.967 | 0.329 |
| Cross-attention maps | 0.966 | 0.329 |
[…]
[table]
Impact of saliency mask
We compare OCD by using saliency masks from three sources: i) human-labeled masks, ii) predictions from a state-of-the-art segmentation models (Grounded-SAM [36]), and iii) cross-attention maps w.r.t. text prompts (available within the model itself). We did not observe meaningful differences either in generations (see Fig. 8 for an example) or in aggregated fidelity metrics (Fig. 8), suggesting that OCD is robust to the source of saliency.
@@
The latency impact of Object-Centric Sampling depends on the size of the foreground areas: in the presence of very small edits it can allow significant savings, whereas for prominent objects its editing cost would be only mildly-affected. We study this behavior in Tab. 2, where we divide 18 sequence-prompt pairs in three sets of 6 pairs each, depending on the number of foreground pixels (or foreground ratio ): Large (), Medium () and Small (). As we show in Tab. 2, in the absence of Object-Centric Sampling, the latency of video editing does not depend on the size of the foreground objects. However, although we remark that our proposal reduces latency for all object sizes, we observe smaller gains on large objects, as the foreground area covers most of the frame resolution, matching Object-Centric Sampling with a regular diffusion process operating on the whole frame. Savings gradually improve as we go towards medium and small sized objects, showing additional speed-ups of 1.3 and 1.8 respectively.
| Object | Avg. #pixel | w/o Obj-Cen. Sampling | w/ Obj-Cen. Sampling | ||||
@@
| size | (fg. ratio ) | Inversion | Generation | Inversion | Generation | ||
| Large | 53.5k (20.4%) | 0.976 | 12.70 | 11.08 | 0.967 | 9.77 (2.93) | 9.65 (1.43) |
| Medium | 35.2k (13.4%) | 0.956 | 13.09 | 11.08 | 0.953 | 9.97 (3.12) | 8.82 (2.26) |
| Small | 17.7k (6.7%) | 0.953 | 13.02 | 10.77 | 0.948 | 7.25 (5.77) | 6.22 (4.55) |
@@ 6 Conclusion
In this paper, we introduced solutions for speeding up diffusion-based video editing. In this respect, we first presented an analysis of sources of latency in inversion-based models, and we identified and adopted some off-the-shelf techniques such as fast sampling and Token Merging that, when properly modified for the task, bring significant cost reduction with only slight quality degradation. Furthermore, motivated by the fact that video editing typically requires modifications to specific objects, we introduce Object-Centric Diffusion, comprising techniques for i) encouraging the merging of tokens in background regions and ii) limiting most of the diffusion sampling steps on foreground areas. Our solutions, which fix generation artifacts on foreground objects and further reduce editing latency, are validated on inversion-based and ControlNet-based models, by achieving 10 and 6 faster edits for comparable quality.
@@ A.1 Off-the-shelf optimizations of ToMe
Many inversion-based image/video editing pipelines rely on sharing attention maps between inversion and generation stages (e.g. FateZero [32], Plug-and-Play [43]). As such, when applying ToMe [3, 4], it is important that locations of destination (dst) and unmerged (unm) tokens are paired in the two stages, at each corresponding attention layer and diffusion step. If that is not the case, tokens or attention maps coming from inversion are not compatible with the ones available at generation time. In practice, we compute which tokens to be merged during inversion, and merge the tokens at the same locations in generation attention maps. By doing so, we make sure the tokens that remain after merging correspond to the same locations (or, token indices), and hence, the attention maps from inversion and generation can rightly be fused. We found this strategy to be of primary importance, as testified by Fig. 3 (d-e) in the main paper.
@@
ToMe for stable diffusion [4] samples dst tokens randomly in a single image. When extending this to multiple frames, we initially sample the same random locations in each frame, finding this strategy to be sub-optimal. Instead, if we re-sample different random locations in each frame (or, in each temporal window in our spatio-temporal implementation), it allows us to preserve different information in each frame (or, window) after merging. We found this to be beneficial, especially at higher merging rates (e.g. see Fig. 3 (e to f) in the main paper).
@@
In the original ToMe for stable diffusion [4], for each source (src) token, we search its corresponding match within a pool of dst tokens coming from the full spatial extent of the given image (). The naive transposition of this strategy to our video use-case allows, for any src token, its candidate match to be searched within a pool of dst tokens coming from the full spatio-temporal extent of the video (). We find that this strategy for searching dst match, named hereby Temporally-Global Search, can lead to generation artifacts. Differently, we consider restricting the temporal extent of the dst pool to be the same temporal-window () as the src token, as in our Temporally-Windowed Search. As shown in Fig. A.1, the latter gives better reconstructions in general, whilst allowing more flexibility to control where merges are happening, temporally. This way, the user can also better trade-off the temporal-redundancy, smoothness and consistency by controlling the spatio-temporal window size.
@@
As observed in [4], the high resolution UNet [38] stages are the most expensive in terms of self-attention (or, cross-frame attention) modules, and the ones that can benefit the most by applying ToMe. Contrarily to the original formulation which does not optimize low-resolution layers, we do apply ToMe in all layers as we observe it has a meaningful impact on latency. We however cap the minimum #tokens preserved after merging in low-resolution layers, in order to avoid degenerate bottlenecks (e.g. collapsing to a single representation). Specifically, we maintain at-least 4 and 16 tokens per-frame after merging at and resolutions, respectively.